OpenStudy (anonymous):

What is the equation of the line perpendicular to y = -x + 3 through (-1,-2)

OpenStudy (anonymous):

if a line has slope m, the line perpendicular to it will have a slope of -1/m (the negative reciprocal)

OpenStudy (anonymous):

you can use point slope form to find the equation (y-y1)=m(x-x1)

OpenStudy (anonymous):

So I re-write it as y = 1/m + 3 first, then write -2+2(1) = 1/x(-1+1(1))?

OpenStudy (anonymous):

to use point slope for you need two things. One point, which you were given, and a slope. You know the new slope (which is usually called "m") you know how to get the slope of a perpendicular line by flipping the fraction upside down (taking the reciprocal) and flipping the sign(from positive to negative or negative to positive)

OpenStudy (anonymous):

If I replaced all the x's, y's and the new slope, how could I solve for anything though? Sorry, this is the first time I've done this type of problem

OpenStudy (ranga):

Let me walk you through step by step: What is the slope of y = -x + 3?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Sorry, meant -1

OpenStudy (ranga):

ok -1 is correct.

OpenStudy (anonymous):

Wasn't thinking, sorry. so the new slope would be 1/x? that's the inverse reciprocal?

OpenStudy (ranga):

The slope of the perpendicular line will be the negative reciprocal of the previous slope: that is -1/(previous slope) So what is the slope of the perpendicular line?

OpenStudy (anonymous):

-1/-1 = 1

OpenStudy (ranga):

correct.

OpenStudy (ranga):

The equation of a line is y = mx + b You know the slope m to be 1. So y = x + b This line passes through the point (-1, -2). So put x = -1 and y = -2 in the above equation and solve for b. b = ?

OpenStudy (anonymous):

b = -1

OpenStudy (ranga):

Yes. Put b = -1 in the equation y = x + b and you now have the equation of the perpendicular line.

OpenStudy (anonymous):

So the answer would be y = 1x -1

OpenStudy (ranga):

Yes. But no need to put the 1 in front of x because 1 multiplied by any number is the same number. So 1x = x.

OpenStudy (anonymous):

Thank you so much, this has been very helpful! Really quick, I had another one of these problems with a perpendicular line and wanted to check if it was correct. It was perp to 5x + 3y = 1 and the answer was y = 3/5x +5

OpenStudy (ranga):

You have to give me the point the perpendicular passes through.

OpenStudy (anonymous):

Oh, I'm sorry. It was (-5,2)

OpenStudy (ranga):

y = 3/5x + 5 You have the correct answer!

OpenStudy (anonymous):

Thank you! I will use your method to solve future problems like this one, the other one confused me a lot. But thank you again for your help!

OpenStudy (ranga):

You are very welcome!